1) The going rate of student loans is quoted as 8 percent APR. The terms of the loans call for monthly payments. What is the effective annual rate (EAR) on such a student loan?
Answer #1

EAR = (1-6)*-1 i = stated annual interest rate n = number of compounding periods

n = 12 (since there are 12 months in a year)

0.08 EAR = (1+ (1 A1 = + )12 -

EAR = 1.083 - 1

EAR = 0.0830

EAR = 8.30%